# Communication-Aware Dissipative Control for Networks of Heterogeneous Nonlinear Agents

### v1

@@ 1 Introduction
Energy-related analysis is widely accepted in the study of stability for physical and engineering systems. Dissipativity (Willems, 1972; Hill and Moylan, 2003) provides an abstract, energy-based characterization of general nonlinear systems by treating them as input-output mappings, ignoring internal states. A key advantage of this abstraction is its design compositionality (Zakeri and Antsaklis, 2022), meaning that the interconnections of dissipative systems are themselves dissipative and stable with minimum conditions. The NDT in (Moylan and Hill, 2003; Vidyasagar, 1981) exploits this to assess network-level stability using only the open-loop properties of individual agents. While the theorem has been successfully applied to the synthesis of decentralized controllers for large-scale networks (LoCicero and Bridgeman, 2025), additional mechanisms are required when communication among controllers is necessary to improve the performance of the controller.
@@ 2.2 $QSR$-Dissipativity of Large-Scale Systems
This section reviews the $QSR$-dissipativity and its application to for stability analysis of networks.
@@
We identify $QSR$-dissipativity of linear time-invariant (LTI) systems as follows.
@@
Consider $N$ dissipative operators, $\mathscr{G}_{i}:\mathcal{L}_{2e}^{m_{i}}\mapsto\mathcal{L}_{2e}^{l_{i}}$ with parameters $(\mathbf{Q}_{i}\mathbf{S}_{i}\mathbf{R}_{i})$, interconnected by the operators, $\mathscr{H}_{ij}{:}\mathcal{L}_{2e}^{l_{j}}{\mapsto}\mathcal{L}_{2e}^{m_{i}}$, mapping the output of the $j^{\mathrm{th}}$ operator to the input of $i^{\mathrm{th}}$ operator and represented by matrices $\mathbf{H}_{ij}{\in}\mathbb{R}^{m_{i}\times l_{j}}$. Thereby, the large-scale interconnected system is expressed by
$\mathbf{y}_{i}=\mathscr{G}\mathbf{e}_{i},\quad\mathbf{y}=\mathscr{G}\mathbf{u},\quad\mathbf{e}=\mathbf{u}+\mathbf{H}\mathbf{y}$ (4)
@@ 2.3 Convex Overbounding
Optimal controller synthesis minimizes an objective function subject to matrix inequality constraints that are typically bilinear matrix inequality (BMI), expressed as
@@
where $\mathbf{N}\in\mathbb{R}^{p\times q}$ is a given matrix, and $\mathbf{Q}\in\mathbb{S}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$. and $\mathbf{Y}\in\mathbb{R}^{q\times n}$ are design variables. Since Equation 6 is generally nonconvex, the associated optimal control problem becomes NP-hard. Conservative convex relaxations of Equation 6 can be derived using the following results.
@@
Directly applying Theorem 3 may introduce excessive conservatism, making it difficult to obtain a feasible solution to Equation 6. Fortunately, its conservatism can be significantly reduced by linearizing around a feasible point, $\mathbf{X}^{0}$ and $\mathbf{Y}^{0}$ satisfying $\mathbf{Q}{+}\text{He}(\mathbf{X}^{0}\mathbf{N}\mathbf{Y}^{0}){\prec}0$, if one is known. Specifically, Theorem 3 can be reformulated by substituting $\mathbf{Q}$, $\mathbf{X}$, and $\mathbf{Y}$ with $\mathbf{Q}{+}\text{He}(\mathbf{X}^{0}\mathbf{N}\mathbf{Y}^{0}{+}\delta{\mathbf{X}}\mathbf{N}\mathbf{Y}^{0}{+}\mathbf{X}^{0}\mathbf{N}\delta{\mathbf{Y}})$, $\delta{\mathbf{X}}$, and $\delta{\mathbf{Y}}$, and solve Equation 7 or 8 with new design variables $\mathbf{Q}$, $\delta{\mathbf{X}}$, and $\delta{\mathbf{Y}}$. The resulting problem is always feasible since $\delta{\mathbf{X}}{=}\mathbf{0}$ and $\delta{\mathbf{Y}}{=}\mathbf{0}$ yields the initial feasible point $(\mathbf{X}^{0},\mathbf{Y}^{0})$.
@@ 3 Sparsity-Promoting Dissipativity-Augmented Control
$\mathbf{e}_{i}$ ${=}\hskip-10.0pt\sum_{j\in\mathbb{N}_{N},i\neq j}\hskip-10.0pt\mathbf{H}_{ij}^{p}\mathbf{x}_{j}{+}\hskip-5.0pt\sum_{j\in\mathbb{N}_{N}}\hskip-3.0pt\widetilde{\mathbf{H}}_{ij}^{p}\mathbf{u}_{j},\quad\widehat{\mathbf{e}}_{i}{=}\hskip-5.0pt\sum_{j\in\mathbb{N}_{N}}\hskip-3.0pt\mathbf{K}_{ij}\widehat{\mathbf{u}}_{ij},$
$\begin{bmatrix}\mathbf{e}\\ \widehat{\mathbf{e}}\end{bmatrix}$ ${=}\begin{bmatrix}\mathbf{H}_{p}&\widetilde{\mathbf{H}}_{p}\\ \mathbf{K}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{x}\\ \widehat{\mathbf{y}}\end{bmatrix}{=}\mathbf{H}\begin{bmatrix}\mathbf{x}\\ \widehat{\mathbf{y}}\end{bmatrix}$,
where $\mathbf{e}=\text{col}(\mathbf{e}_{i})_{i\in\mathbb{N}_{N}}$, and $\widehat{\mathbf{e}}$, $\mathbf{x}$, and $\widehat{\mathbf{y}}$ are defined analogously. $\mathbf{H}_{p}$ characterizes the interconnection structure among agent, $\widetilde{\mathbf{H}}_{p}$ represents the interconnection structure from the input of the plant to the input of each agent, and $\mathbf{H}_{ij}^{p}$ and $\widetilde{\mathbf{H}}_{ij}^{p}$ are the corresponding block components. Figure 1 depicts this system decomposition. The objective is to synthesize a sparse controller $\mathbf{K}$ that stabilizes the overall interconnected system in the presence of disturbances $\mathbf{r}$ and $\widehat{\mathbf{r}}$, while minimizing a standard LQ criterion.
@@ 3.1 Dual-Model Synthesis
Stability is ensured by enforcing $QSR$-dissipativity properties on each agent and its local controller, $\mathcal{G}_{i}$ and $\mathcal{C}_{i}$, and by applying Theorem 2. Although the external inputs to the decomposed network are $\widetilde{\mathbf{H}}_{p}\mathbf{r}$ and $\mathbf{K}\widehat{\mathbf{r}}$, Theorem 2 remains applicable because signals remain in $\mathcal{L}_{2}$ after premultiplication by a constant gain.
The LQ criterion is evaluated based on the closed-loop dynamics of the linearized plant, $\mathcal{G}^{LTI}$, and controller, described as
@@
where $\mathbf{A}$ and $\mathbf{B}$ denote the linearized system matrices of the plant. Synthesizing $\mathbf{K}$ matrix that minimizes LQ criterion of the closed-loop system is equivalent to solving
@@
where $\mathbf{Q}_{lq}\succ 0$ and $\mathbf{R}_{lq}\succ 0$ are state and control weight of the LQ criterion (Khalil et al., 1996). To promote the sparsity in $\mathbf{K}$, a penalty function $g(\mathbf{K})$ is incorporated into the objective function. Typical sparsity-promoting penalty functions include the $\ell_{1}$ norm, weighted $\ell_{1}$ norm, sum-of-logs, and the cardinality function of $\mathbf{K}$ (Lin et al., 2013).
@@
where $\gamma$ denotes a penalty parameter, $(\mathbf{Q}_{i},\mathbf{S}_{i},\mathbf{R}_{i})$ and $(\widehat{\mathbf{Q}}_{i},\widehat{\mathbf{S}}_{i},\widehat{\mathbf{R}}_{i})$ respectively represents $QSR$ parameters of $i^{\mathrm{th}}$ agent and its local controller, and $\mathbf{Q}=\text{diag}(\text{diag}(\mathbf{Q}_{i})_{i\in\mathbb{N}_{N}},\text{diag}(\widehat{\mathbf{Q}}_{i})_{i\in\mathbb{N}_{N}})$, while $\mathbf{S}$ and $\mathbf{R}$ are defined analogously. However, Equation 9 is non-convex optimization problem, due to Equation 9b, 9c and 9e. These constraints are relaxed to linear matrix inequalities in Section 4.
@@ 4 Main Results
In this section, the strategies solving Equation 9 are proposed. A convex relaxation of Equation 9 is presented to obtain a feasible solution in a computationally tractable manner. Equation 9c is reformulated to linear matrix inequalitys (LMIs) by applying the expansions from Lemma 1, based on the structures of agents. Equation 9b and 9e are reconstructed using the convex overbounding techniques introduced in Section 2.3. Section 4.4 provides the approach to compute a feasible sparse controller than improves the LQ performance criterion, based on ADMM.
@@ 4.1 Agent Dissipativity
Definition 1 cannot be directly applied to solve Equation 9 using the existing semidefinite programming (SDP)-based optimization solvers. Fortunately, many variations of Definition 1 result in tractable LMIs that enforce the $QSR$-dissipativity of certain nonlinear systems. For agents $\mathcal{G}_{i}$ modeled as LTI systems, Lemma 1 can be applied to impose the dissipativity constraint. If $\mathcal{G}_{i}$ is an LTI system with polytopic uncertainty in its dynamics, (Walsh and Forbes, 2019, Equation 18) can be used to verify the $QSR$-dissipativity of the agent. In case where $\mathcal{G}_{i}$ an LTI system with time delays in its input, output, and state, the results in (Bridgeman and Forbes, 2016, Theorem 3.1) provides an LMI condition for determining dissipativity parameters of the agent.
If no existing LMI condition ensures the dissipativity of $\mathcal{G}_{i}$, the predefined dissipativity parameters $(\mathbf{Q}_{i}^{p},\mathbf{S}_{i}^{p},\mathbf{R}_{i}^{p})$ can be used together with a scalar design variable $\lambda$, such that $\mathbf{Q}_{i}=\lambda\mathbf{Q}_{i}^{p}$, $\mathbf{S}_{i}=\lambda\mathbf{S}_{i}^{p}$, and $\mathbf{R}_{i}=\lambda\mathbf{R}_{i}^{p}$. Although these parameters are fixed, introducing the scalar variable $\lambda$ alleviates the conservatism inherent in the fixed characterizations. Various approaches exist to find the dissipativity of nonlinear agents. For instance, agents following Eulter-Lagrange systems is passive, meaning $\mathbf{Q}_{i}^{p}=\mathbf{0}$, $\mathbf{S}_{i}^{p}=\frac{1}{2}\mathbf{I}$, and $\mathbf{R}_{i}^{p}=\mathbf{0}$ (Lozano et al., 2013, Chapter 6). In addition, data-driven methods such as (Romer et al., 2017) can be employed to estimate $(\mathbf{Q}_{i}^{p},\mathbf{S}_{i}^{p},\mathbf{R}_{i}^{p})$ for the agents.
@@ 4.2 Convex Overbounding of BMI Constraints
Convexification tricks in Theorem 3 can be applied to yield LMIs implying Equation 9b and 9e, explained in Corollary 1 and 2.
@@
Suppose that $\mathbf{K}^{0}$ and $\mathbf{P}^{0}$ are feasible points of Equation 9b. For any $\mathbf{G}^{0}$ and $\mathbf{F}^{0}$ satisfying $\text{He}(\mathbf{G}^{0})\succeq 0$ and $\mathbf{F}^{0}\succeq 0$,
@@
where $\mathbf{L}{=}\text{He}\{\hskip-1.5pt(\mathbf{P}^{0}{+}\delta{\mathbf{P}})(\mathbf{A}{-}\mathbf{B}\mathbf{K}^{0}){-}\delta{\mathbf{P}}\mathbf{B}\mathbf{K}^{0}\hskip-1.5pt\}$, implies that $\mathbf{K}^{0}{+}\delta{\mathbf{K}}$ and $\mathbf{P}^{0}{+}\delta{\mathbf{P}}$ are feasible solution of Equation 9b.
@@
The proof follows by applying Schur complement and the overbounding condition in Equation 8 of Theorem 3. Since $\mathbf{R}_{lq}\succ 0$, applying Schur complement to Equation 9b and substituting $\mathbf{K}$ and $\mathbf{P}$ with $\mathbf{K}^{0}{+}\delta{\mathbf{K}}$ and $\mathbf{P}^{0}{+}\delta{\mathbf{P}}$ yield
@@
Suppose that $\mathbf{Q}^{0}$, $\mathbf{S}^{0}$, $\mathbf{R}^{0}$, and $\mathbf{H}^{0}$ are feasible points of Equation 9e. Then
$\begin{bmatrix}\mathbf{M}&*&*&*\\ \delta{\mathbf{S}}^{T}{+}\delta{\mathbf{H}}&{-}2\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \frac{1}{2}\mathbf{R}^{0}\delta{\mathbf{H}}{+}\delta{\mathbf{H}}&\mathbf{0}&{-}2\mathbf{I}&*\\ \delta{\mathbf{R}}\mathbf{H}^{0}{+}\delta{\mathbf{H}}&\mathbf{0}&\frac{1}{2}\delta{\mathbf{R}}&{-}2\mathbf{I}\end{bmatrix}{\prec}0$ (11)
where $\mathbf{M}{=}\mathbf{Q}^{0}\hskip-1.0pt{+}\delta{\mathbf{Q}}\hskip-1.0pt{+}\hskip-1.0pt\text{He}(\hskip-1.0pt\mathbf{S}^{0}\mathbf{H}^{0}{+}\delta{\mathbf{S}}\mathbf{H}^{0}{+}\mathbf{S}^{0}\delta{\mathbf{H}}\hskip-1.0pt){+}\mathbf{H}^{0T}\hskip-1.0pt\mathbf{R}^{0}\mathbf{H}^{0}\hskip-5.0pt+\mathbf{H}^{0T}\delta{\mathbf{R}}\mathbf{H}^{0}{+}\text{He}(\delta{\mathbf{H}}^{T}\mathbf{R}^{0}\mathbf{H}^{0})$, implies that $\mathbf{Q}^{0}{+}\delta{\mathbf{Q}}$, $\mathbf{S}^{0}{+}\delta{\mathbf{S}}$, $\mathbf{R}^{0}{+}\delta{\mathbf{R}}$, and $\mathbf{H}^{0}{+}\delta{\mathbf{H}}$ are feasible solution of Equation 9e.
@@
Since Equation 9e is a trilinear matrix inequality and $\mathbf{R}\nsucc 0$, Equation 11 is obtained by applying Equation 7 of Theorem 3 twice to Equation 9e. First, using $\mathbf{S}=\mathbf{S}^{0}+\delta{\mathbf{S}}$ and $\mathbf{H}=\mathbf{H}^{0}+\delta{\mathbf{H}}$ allows Equation 9e to be reformulated as
@@
$\quad{+}\text{He}\left(\begin{bmatrix}\delta{\mathbf{S}}&\frac{1}{2}\delta{\mathbf{H}}^{T}\mathbf{R}\end{bmatrix}\begin{bmatrix}\delta{\mathbf{H}}\\ \delta{\mathbf{H}}\end{bmatrix}\right){\prec}0$
Applying Equation 7 of Theorem 3 with $\mathbf{X}=[\delta{\mathbf{S}}\;\frac{1}{2}\delta{\mathbf{H}}^{T}\mathbf{R}]$, $\mathbf{Y}{=}[\delta{\mathbf{H}}^{T}\;\delta{\mathbf{H}}^{T}]^{T}$, $\mathbf{N}{=}\mathbf{I}$ and $\mathbf{G}{=}\mathbf{I}$, and replacing $\mathbf{Q}$ and $\mathbf{R}$ with $\mathbf{Q}^{0}{+}\delta{\mathbf{Q}}$ and $\mathbf{R}^{0}{+}\delta{\mathbf{R}}$, respectively, lead to
@@
and $\delta{\mathbf{S}}$ and $\delta{\mathbf{R}}$ are defined likewise. Applying Corollary 1 and 2, we have the problem
@@
to construct the centralized controller for given initial feasible point, where $\mathbf{N}(\delta{\mathbf{Q}},\delta{\mathbf{S}},\delta{\mathbf{R}},\delta{\mathbf{K}},\delta{\mathbf{P}},\delta{\mathbf{G}},\mathbf{F})$ denotes the block diagonal composition of all matrices associated with the LMIs.

### v2

@@ 1 Introduction
Energy-related analysis is widely accepted in the study of stability for physical and engineering systems. Dissipativity (Willems, 1972; Hill and Moylan, 2003) provides an abstract, energy-based characterization of general nonlinear systems by treating them as input-output mappings, ignoring internal states. A key advantage of this abstraction is its design compositionality (Zakeri and Antsaklis, 2022), meaning that the interconnections of dissipative systems are themselves dissipative and stable with minimum conditions. The NDT in (Moylan and Hill, 2003; Vidyasagar, 1981) exploits this to assess network-level stability using only the open-loop properties of individual agents. While the theorem has been successfully applied to the synthesis of decentralized controllers for large-scale networks (Arcak et al., 2016; LoCicero and Bridgeman, 2025), additional mechanisms are required when communication among controllers is necessary to improve the performance of the controller.
@@ 2.2 $QSR$-Dissipativity of Large-Scale Systems
This section reviews the $QSR$-dissipativity and its application to stability analysis of networks.
@@
We identify $QSR$-dissipativity of linear time-invariant (LTI) systems using the following lemma. Note that the stability assumption in (Gupta, 1996) was only needed for necessity. Hence, it provides sufficient conditions for dissipativity of both stable and unstable systems.
@@
Consider $N$ dissipative operators, $\mathscr{G}_{i}{:}\mathcal{L}_{2e}^{m_{i}}{\mapsto}\mathcal{L}_{2e}^{l_{i}}$ with parameters $(\mathbf{Q}_{i}\mathbf{S}_{i}\mathbf{R}_{i})$, interconnected by the operators, $\mathscr{H}_{ij}:\mathcal{L}_{2e}^{l_{j}}{\mapsto}\mathcal{L}_{2e}^{m_{i}}$, mapping the output of $j^{\mathrm{th}}$ operator to the input of $i^{\mathrm{th}}$ operator and represented by matrices $\mathbf{H}_{ij}{\in}\mathbb{R}^{m_{i}\times l_{j}}$. Thereby, the large-scale interconnected system is expressed by
$\mathbf{y}_{i}=\mathscr{G}\mathbf{e}_{i},\quad\mathbf{y}=\mathscr{G}\mathbf{u},\quad\mathbf{e}=\mathbf{u}+\mathbf{H}\mathbf{y},$ (4)
@@ 2.3 Convex Overbounding
Optimal controller synthesis minimizes an objective function subject to matrix inequality constraints that are typically bilinear matrix inequalities (BMIs), expressed as
@@
where $\mathbf{N}{\in}\mathbb{R}^{p{\times}q}$ is a given matrix, and $\mathbf{Q}{\in}\mathbb{S}^{n}$, $\mathbf{X}{\in}\mathbb{R}^{n{\times}p}$. and $\mathbf{Y}{\in}\mathbb{R}^{q{\times}n}$ are design variables. Since ()6 is generally nonconvex, the associated optimal control problem becomes NP-hard. Conservative convexifications of ()6 can be derived using the following results.
@@
Directly applying Theorem 3 may introduce excessive conservatism, making it difficult to obtain a feasible solution to ()6. Fortunately, its conservatism can be significantly reduced by linearizing around a feasible point, $\mathbf{X}^{0}$ and $\mathbf{Y}^{0}$ satisfying $\mathbf{Q}{+}\text{He}(\mathbf{X}^{0}\mathbf{N}\mathbf{Y}^{0}){\prec}0$, if one is known. Specifically, Theorem 3 can be reformulated by substituting $\mathbf{Q}$, $\mathbf{X}$, and $\mathbf{Y}$ with $\mathbf{Q}{+}\text{He}(\mathbf{X}^{0}\mathbf{N}\mathbf{Y}^{0}{+}\delta{\mathbf{X}}\mathbf{N}\mathbf{Y}^{0}{+}\mathbf{X}^{0}\mathbf{N}\delta{\mathbf{Y}})$, $\delta{\mathbf{X}}$, and $\delta{\mathbf{Y}}$, and solve ()7 or ()8 with new design variables $\mathbf{Q}$, $\delta{\mathbf{X}}$, and $\delta{\mathbf{Y}}$. The resulting problem is always feasible since $\delta{\mathbf{X}}{=}\mathbf{0}$ and $\delta{\mathbf{Y}}{=}\mathbf{0}$ yields the initial feasible point $(\mathbf{X}^{0},\mathbf{Y}^{0})$.
@@ 3 Sparsity-Promoting Dissipativity-Augmented Control
$\mathbf{e}_{i}$ ${=}\hskip-10.0pt\sum_{j\in\mathbb{N}_{N},i\neq j}\hskip-10.0pt\mathbf{H}_{ij}^{p}\mathbf{x}_{j}{+}\hskip-5.0pt\sum_{j\in\mathbb{N}_{N}}\hskip-3.0pt\widetilde{\mathbf{H}}_{ij}^{p}\mathbf{u}_{j},\quad\widehat{\mathbf{e}}_{i}{=}\hskip-5.0pt\sum_{j\in\mathbb{N}_{N}}\hskip-3.0pt\mathbf{K}_{ij}\widehat{\mathbf{u}}_{j},$
$\begin{bmatrix}\mathbf{e}\\ \widehat{\mathbf{e}}\end{bmatrix}$ ${=}\begin{bmatrix}\widetilde{\mathbf{H}}_{p}\mathbf{r}\\ \mathbf{K}\widehat{\mathbf{r}}\end{bmatrix}+\begin{bmatrix}\mathbf{H}_{p}&\widetilde{\mathbf{H}}_{p}\\ \mathbf{K}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{x}\\ \widehat{\mathbf{y}}\end{bmatrix}{=}\begin{bmatrix}\widetilde{\mathbf{H}}_{p}\mathbf{r}\\ \mathbf{K}\widehat{\mathbf{r}}\end{bmatrix}+\mathbf{H}\begin{bmatrix}\mathbf{x}\\ \widehat{\mathbf{y}}\end{bmatrix}$,
where $\mathbf{e}=\text{col}(\mathbf{e}_{i})_{i\in\mathbb{N}_{N}}$, and $\widehat{\mathbf{e}}$, $\mathbf{x}$, and $\widehat{\mathbf{y}}$ are defined analogously. $\mathbf{H}_{p}$ characterizes the interconnection structure among agents, $\widetilde{\mathbf{H}}_{p}$ represents the interconnection structure from the input of the plant to the input of each agent, and $\mathbf{H}_{ij}^{p}$ and $\widetilde{\mathbf{H}}_{ij}^{p}$ are the corresponding block components. Figure 1 depicts this system decomposition. The objective is to synthesize a sparse controller $\mathbf{K}$ that stabilizes the overall interconnected system in the presence of disturbances $\mathbf{r}$ and $\widehat{\mathbf{r}}$, while minimizing a standard LQ criterion.
@@ 3.1 Dual-Model Synthesis
Stability is ensured by enforcing $QSR$-dissipativity properties on each agent and its local controller, $\mathcal{G}_{i}$ and $\mathcal{C}_{i}$, and by applying NDT, Theorem 2. Although the external inputs to the decomposed network are $\widetilde{\mathbf{H}}_{p}\mathbf{r}$ and $\mathbf{K}\widehat{\mathbf{r}}$, Theorem 2 remains applicable because signals remain in $\mathcal{L}_{2}$ after premultiplication by a constant gain.
The LQ performance is evaluated based on the closed-loop dynamics of the linearized plant, $\mathcal{G}^{LTI}$, and controller,
@@
where $\mathbf{A}$ and $\mathbf{B}$ denote the linearized plant matrices. Synthesizing $\mathbf{K}$ matrix that minimizes LQ cost of the closed-loop system is equivalent to solving
@@
where $\mathbf{Q}_{lq}\succ 0$ and $\mathbf{R}_{lq}\succ 0$ are state and control weight of the LQ criterion (Khalil et al., 1996). To promote sparsity in $\mathbf{K}$, a penalty of $g(\mathbf{K})$ is added to the objective. Typical sparsity-promoting penalty functions include the $\ell_{1}$ norm, weighted $\ell_{1}$ norm, sum-of-logs, and the cardinality function of $\mathbf{K}$ (Lin et al., 2013).
@@
where $\gamma$ denotes a penalty parameter, $(\mathbf{Q}_{i},\mathbf{S}_{i},\mathbf{R}_{i})$ and $(\widehat{\mathbf{Q}}_{i},\widehat{\mathbf{S}}_{i},\widehat{\mathbf{R}}_{i})$ respectively represent $QSR$ parameters of $i^{\mathrm{th}}$ agent and its local controller, and $\mathbf{Q}=\text{diag}(\text{diag}(\mathbf{Q}_{i})_{i\in\mathbb{N}_{N}},\text{diag}(\widehat{\mathbf{Q}}_{i})_{i\in\mathbb{N}_{N}})$, while $\mathbf{S}$ and $\mathbf{R}$ are defined analogously. However, ()9 is non-convex due to ()9b, ()9c, and ()9e. These constraints are tightened to linear matrix inequalities (LMIs) in Section 4.
@@ 4 Main Results
This section proposes strategies to solve ()9. ()9c is reformulated to LMIs by applying the expansions from Lemma 1, based on the structures of agents. ()9b, and ()9e are reconstructed using the convex overbounding techniques introduced in Section 2.3. Section 4.4 provides the approach to compute a feasible sparse controller that improves the LQ performance criterion, based on ADMM.
@@ 4.1 Agent Dissipativity
Definition 1 cannot be directly applied to solve ()9 using existing semidefinite programming (SDP)-based optimization solvers. Fortunately, many variations of Definition 1 result in tractable LMIs that enforce the $QSR$-dissipativity of certain nonlinear systems. For agents $\mathcal{G}_{i}$ modeled as LTI systems, Lemma 1 can be applied to impose the dissipativity constraint. If $\mathcal{G}_{i}$ is an LTI system with polytopic uncertainty in its dynamics, (Walsh and Forbes, 2019, Equation 18) can be used to verify the $QSR$-dissipativity of the agent. If $\mathcal{G}_{i}$ is an LTI system with input, output or state delays, (Bridgeman and Forbes, 2016, Theorem 3.1) provides LMIs to determine dissipativity parameters of the agent. Gain of control affine systems can be determined from LMIs following Strong et al. (2024).
In practice, dissipativity of nonlinear agents is often established a priori through ad hoc analysis, or prior knowledge. For instance, Eulter-Lagrange systems are passive, meaning $\mathbf{Q}_{i}^{p}=\mathbf{0}$, $\mathbf{S}_{i}^{p}=\frac{1}{2}\mathbf{I}$, and $\mathbf{R}_{i}^{p}=\mathbf{0}$ (Lozano et al., 2013, Chapter 6). In addition, data-driven methods such as (Romer et al., 2017) can be employed to estimate $(\mathbf{Q}_{i}^{p},\mathbf{S}_{i}^{p},\mathbf{R}_{i}^{p})$ for the agents. Then the predefined dissipativity parameters $(\mathbf{Q}_{i}^{p},\mathbf{S}_{i}^{p},\mathbf{R}_{i}^{p})$ can be scaled by a design variable $\lambda{\geq}0$, so that $\mathbf{Q}_{i}{=}\lambda\mathbf{Q}_{i}^{p}$, $\mathbf{S}_{i}{=}\lambda\mathbf{S}_{i}^{p}$, and $\mathbf{R}_{i}{=}\lambda\mathbf{R}_{i}^{p}$ to search over equivalent characterizations.
@@ 4.2 Convex Overbounding of BMI Constraints
Corollaries 1 and 2 apply Theorem 3 to yield LMIs implying ()9b, and ()9e.
@@
Given $\mathbf{K}^{0}$ and $\mathbf{P}^{0}$, suppose there exist $\delta{\mathbf{K}}$, $\delta{\mathbf{P}}$, $\mathbf{G}^{0}$ and $\mathbf{F}^{0}{\succeq}0$ satisfying $\text{He}(\mathbf{G}^{0}){\succeq}0$ and
@@
where $\mathbf{L}{=}\text{He}\{\hskip-1.5pt(\mathbf{P}^{0}{+}\delta{\mathbf{P}})(\mathbf{A}{-}\mathbf{B}\mathbf{K}^{0}){-}\mathbf{P}^{0}\mathbf{B}\delta{\mathbf{K}}\}\hskip-1.5pt$. Then $\mathbf{K}^{0}+\delta{\mathbf{K}}$ and $\mathbf{P}^{0}{+}\delta{\mathbf{P}}$ are feasible points of ()9b. Moreover, if $\mathbf{K}^{0}$ and $\mathbf{P}^{0}$ are feasible for ()9b, ()10 is always feasible.
@@
The proof follows by applying Schur complement and the overbounding condition in ()8 of Theorem 3. Since $\mathbf{R}_{lq}{\succ}0$, applying Schur complement to ()9b and substituting $\mathbf{K}$ and $\mathbf{P}$ with $\mathbf{K}^{0}{+}\delta{\mathbf{K}}$ and $\mathbf{P}^{0}{+}\delta{\mathbf{P}}$ yield
@@
Given $\mathbf{Q}^{0}$, $\mathbf{S}^{0}$, $\mathbf{R}^{0}$, and $\mathbf{H}^{0}$, suppose there exists $\delta{\mathbf{Q}}$, $\delta{\mathbf{S}}$, $\delta{\mathbf{R}}$, and $\delta{\mathbf{H}}$ such that
$\begin{bmatrix}\mathbf{M}&*&*&*\\ \delta{\mathbf{S}}^{T}{+}\delta{\mathbf{H}}&{-}2\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \frac{1}{2}\mathbf{R}^{0}\delta{\mathbf{H}}{+}\delta{\mathbf{H}}&\mathbf{0}&{-}2\mathbf{I}&*\\ \delta{\mathbf{R}}\mathbf{H}^{0}{+}\delta{\mathbf{H}}&\mathbf{0}&\frac{1}{2}\delta{\mathbf{R}}&{-}2\mathbf{I}\end{bmatrix}{\prec}0,$ (11)
where $\mathbf{M}{=}\mathbf{Q}^{0}\hskip-1.0pt{+}\delta{\mathbf{Q}}\hskip-1.0pt{+}\hskip-1.0pt\text{He}(\hskip-1.0pt\mathbf{S}^{0}\mathbf{H}^{0}{+}\delta{\mathbf{S}}\mathbf{H}^{0}{+}\mathbf{S}^{0}\delta{\mathbf{H}}\hskip-1.0pt){+}\mathbf{H}^{0T}\hskip-1.5pt\mathbf{R}^{0}\mathbf{H}^{0}\hskip-5.0pt+\mathbf{H}^{0T}\hskip-1.0pt\delta{\mathbf{R}}\mathbf{H}^{0}\hskip-1.5pt{+}\text{He}(\hskip-1.0pt\delta{\mathbf{H}}^{T}\mathbf{R}^{0}\mathbf{H}^{0}\hskip-1.0pt)$. Then $\mathbf{Q}^{0}\hskip-2.0pt{+}\delta{\mathbf{Q}}$, $\mathbf{S}^{0}\hskip-2.0pt{+}\delta{\mathbf{S}}$, $\mathbf{R}^{0}\hskip-2.0pt{+}\delta{\mathbf{R}}$, and $\mathbf{H}^{0}\hskip-2.0pt{+}\delta{\mathbf{H}}$ are a feasible solution of ()9e. Moreover, if $\mathbf{Q}^{0}$, $\mathbf{S}^{0}$, $\mathbf{R}^{0}$, and $\mathbf{H}^{0}$ are feasible for ()9e, ()11 is always feasible.
@@
Since ()9e is a trilinear matrix inequality, ()11 is obtained by applying ()7 of Theorem 3 twice to ()9e. First, using $\mathbf{S}{=}\mathbf{S}^{0}{+}\delta{\mathbf{S}}$ and $\mathbf{H}{=}\mathbf{H}^{0}{+}\delta{\mathbf{H}}$ allows ()9e to be reformulated as
@@
$\quad{+}\text{He}\left(\begin{bmatrix}\delta{\mathbf{S}}&\frac{1}{2}\delta{\mathbf{H}}^{T}\mathbf{R}\end{bmatrix}\begin{bmatrix}\delta{\mathbf{H}}\\ \delta{\mathbf{H}}\end{bmatrix}\right){\prec}0.$
Applying ()7 of Theorem 3 with $\mathbf{X}=[\delta{\mathbf{S}}\;\frac{1}{2}\delta{\mathbf{H}}^{T}\mathbf{R}]$, $\mathbf{Y}{=}[\delta{\mathbf{H}}^{T}\;\delta{\mathbf{H}}^{T}]^{T}$, $\mathbf{N}{=}\mathbf{I}$ and $\mathbf{G}{=}\mathbf{I}$, and replacing $\mathbf{Q}$ and $\mathbf{R}$ with $\mathbf{Q}^{0}{+}\delta{\mathbf{Q}}$ and $\mathbf{R}^{0}{+}\delta{\mathbf{R}}$, respectively, lead to
@@
and $\delta{\mathbf{S}}$ and $\delta{\mathbf{R}}$ are defined like $\delta{\mathbf{Q}}$. Applying Corollary 1 and Corollary 2, we have the problem
@@
to construct the centralized controller for given initial feasible point, where $\mathbf{N}(\delta{\mathbf{Q}},\delta{\mathbf{S}},\delta{\mathbf{R}},\delta{\mathbf{K}},\delta{\mathbf{P}},\delta{\mathbf{G}},\mathbf{F})$ composes: ()10 to ensure that $\text{Trace}(\mathbf{P}^{0}+\delta{\mathbf{P}})$ bounds the LQ norm above; agent dissipativity requirements encoded as LMIs, as discussed in Section 4.1; ()9d to impose controller dissipativity; and ()11 to ensure that the network is stable via NDT.